Flawless Cosmetic Company manufactures and distributes several different products. The company currently uses a plantwide allocation method for allocating overhead at a rate of $7 per direct labor hour. Loren is the department manager of the Makeup Department that produces Products – Concealer (C) and Glow Cream (GC). Jennifer is the department manager of the Hair Care Department that manufactures Product – Shampoo (S). The product costs (per case of 24 bottles) and other information are as follows:

Products
C GC S
Direct materials $ 95.00 $ 76.00 $ 48.00
Direct labor 44.00 34.50 10.00
Overhead 24.00 19.00 15.00
$ 163.00 $ 129.50 $ 73.00
Machine hours 5 3 4
Number of cases (per year) 250 450 550

If Flawless changes its allocation basis to machine hours, what is the total product cost per case for Product GC?
Answer #1

Answer:

Total Overhead = $24 x 250 + $19 x 450 + $15 x 550 = $22800

Total Machine hours = 5 x 250 + 3 x 450 + 4 x 550 = 4800

Overhead Rate = $22800 / 4800 = $4.75 per hour

Product cost per Case for Product GC = 76 + 34.50 + ( $4.75 X 3 ) = $124.75